OpenStudy (anonymous):

1. What is the simplified form of x^4-81/x+3 ? (1 point) a.x3 – 3x2 + 9x – 27 b.(x2 + 9)(x2 – 9) c.x3 + 3x2 + 9x + 27 d.x3 + 3x2 – 9x – 27

OpenStudy (anonymous):

\[\frac{(x^2-9)(x^2+9)}{x+3}=\frac{(x-3)(x+3)(x^2+9)}{x+3}=(x-3)(x^2+9)=x^3-3x^2+9x-27\]